OpenStudy (anonymous):

The solution to the radical function is y = 25.Answer True False

OpenStudy (anonymous):



OpenStudy (anonymous):

@jennychan12

OpenStudy (jennychan12):

ok, then first square both sides. \[3\sqrt{y+2} = 3\] \[9(x+2) = 9\] 9x+18 = 81 solve for x.

OpenStudy (jennychan12):

whoops that y should be an x

OpenStudy (anonymous):

So false! or true?

OpenStudy (anonymous):

@jennychan12 false?

OpenStudy (jennychan12):

what oh, whoops; no. y = 7

OpenStudy (anonymous):

The answer is false! (:

OpenStudy (jennychan12):

yeah sorry.

OpenStudy (anonymous):

it okay! (:

OpenStudy (anonymous):

wait no?

OpenStudy (anonymous):

I'm lost...

OpenStudy (anonymous):

@jennychan12 False yes?

OpenStudy (jennychan12):

yeah false.